OpenStudy (anonymous):

Identify the 16th term of a geometric sequence where a1 = 4 and a8 = -8,748.

OpenStudy (tkhunny):

a1 = 4 a2 = 4r a3 = 4r^2 a4 = 4r^3 ... a8 = 4r^7 = -8748 Can you find r? Do you need to find r? a9 = 4r^8 ... a16 = 4r^15 Okay, maybe you need to find r. The 17th term would have been easier.

OpenStudy (anonymous):

so im confused in what to do next @tkhunny

OpenStudy (tkhunny):

Hmmm... Five hours and no where? a1 = 4 a8 = 4r^7 = -8748 So, r^7 = -2,187 And, t = -2,187^(1/7) = -3 One more thing.